# CardioComposer: Leveraging Differentiable Geometry for Compositional Control of Anatomical Diffusion Models

### v2

@@ Abstract
Generative models of 3D cardiovascular anatomy can synthesize informative structures for clinical research and medical device evaluation, but face a trade-off between geometric controllability and realism. We propose CardioComposer: a programmable, inference-time framework for generating multi-class anatomical label maps based on interpretable ellipsoidal primitives. These primitives represent geometric attributes such as the size, shape, and position of discrete substructures. We specifically develop differentiable measurement functions based on voxel-wise geometric moments, enabling loss-based gradient guidance during diffusion model sampling. We demonstrate that these losses can constrain individual geometric attributes in a disentangled manner and provide compositional control over multiple substructures. Finally, we show that our method is compatible with a wide array of anatomical systems containing non-convex substructures, spanning cardiac, vascular, and skeletal organs.
@@ 1 Introduction
Three-dimensional segmentations of human anatomy power a variety of physics-based simulation platforms. For example, virtual cohorts of anatomy can be used for virtual clinical trials to evaluate and optimize novel medical devices and imaging systems sarrami2021silico; viceconti2021possibleinsilico; abadi2020virtual. Additionally, biophysical simulations can generate insights in the context of both computational physiology research niederer2020creation; straughan2023fully; roney2020silicoablation and surgical training yu2024orbit. Anatomical segmentations can also be used to augment machine-learning workflows through the formation of synthetic images, either through imaging simulators gopalakrishnan2024intraoperative; gopalakrishnan2022fast, domain randomization dey2024learning; billot2023synthseg, or generative models fernandez2022can; fernandez2024generating.
@@
Generative models of anatomy trained on patient-specific data offers advantages for simulation use-cases. For example, conditional generation can augment computational trial cohorts with anatomical variants that are both novel and rare kong2024sdf4chd. Moreover, generative editing methods, such as inpainting, can precisely modify existing patient geometries to create anatomically plausible variations kadry2024probing; kadry2025diffusion. These “digital siblings” can be used with biophysical simulators to model counterfactual scenarios that elucidate the relationship between anatomical form and function.
However, unlike generative modeling of 3D shapes for artistic purposes, generating anatomical models for biophysical simulations presents several unique challenges. The first challenge concerns scale-critical features, in which minor geometric variations on the order of millimeters can cause major fluctuations in physiological behavior fabris2022thinsmall; sacco2018leftsmall; moore2015coronarysmall. Second, anatomical structure exerts attribute-specific effects, in which geometric features such as size and position play different roles in determining biophysical outcomes kadry2021platform. Third, the geometry of multiple substructures interact in a compositional manner kadry2021platform; bhalodia2018deepssm; kong2023sdf4chd, where simulated outcomes are influenced by the collective arrangement of various substructures. Lastly, to interface with clinicians and device engineers, such generative models should be controllable via primitives that are interpretable and physiologically relevant.
@@
  • Differentiable Geometry for Anatomical Characterization: We introduce a set of differentiable geometric measurement functions that compute physiologically relevant anatomical features from a substructure label map. We specifically measure voxel-wise geometric moments compute size via zeroth-order moments, position via first-order moments, and shape via scale-normalized second-order moments.
@@ 2 Background
Traditional Morphometric Modeling for Anatomy. Morphometry involves quantifying anatomical structure through geometric measurements. Traditional morphometric approaches measure intrinsic features such as length, area, volume, and shape, as well as extrinsic factors such as position and orientation. Geometric measurements enable applications such as cardiovascular risk stratification asheghan2023predictingssm; mahmod2023differentiatingstatisticalshape and orthopedic diagnosis gatti2024shapemed. However, traditional morphometric approaches face two key challenges: they cannot represent complex relationships between features, and multiple distinct anatomies may map to the same high-level measurements. To address these limitations, we propose a framework in which an unconditional diffusion model is controlled by traditional morphometric features (size, shape, and position) to generate a variety of realistic anatomical structures, providing an approachable anatomical modeling interface for both clinical and engineering workflows involving numerical simulation.
@@
Generative Models for Numerical Simulation. Anatomical models in the form of 3D meshes or label maps, serve as a crucial tool for studying form-function relationships through physical simulations, enabling both scientific discovery and medical device design. However, current approaches to create such models must trade off between fidelity and control. Simple geometric primitives, such as cylinders for coronary arteries madani2019bridgingfemdeeplearningstress and truncated ellipsoids for cardiac chambers arostica2025software offer parametric control but fail to capture anatomical realism. Data-driven approaches such as autoencoders dou2022generativechimeras; qiao2025personalized represent anatomy in terms of global shape vectors, and can generate synthetic data for mechanistic studies of heart disease hermida2024onsetdeviccistatisticalshape; williams2022aortic. However, such approaches are limited in their ability to model interpretable geometric attributes. Deformation editing methods pham2023svmorph; pham2024virtualshape allow for interpretable control of anatomical geometry, but are limited to modifying existing models. Recently, diffusion-based approaches such as inpainting and partial diffusion have been used to edit patient-specific anatomy to create “digital siblings” kadry2024probing. However, such edits can induce undesirable morphological bias when applied to rare and pathological cases. To this end, recent studies have imposed anatomical features by explicitly providing scalar conditioning features during training. For example, de2025steerable trained a conditional model on thyroid segmentations, and  kadry2025diffusion introduced morpho-skeletal conditioning and guidance mechanisms for coronary arteries. However, both approaches rely on conditional training and are restricted to size-related variables such as volume or cross-sectional area. Similarly, du2025hug presented a hierarchical conditional diffusion model for generating aortic centerlines and radial profiles, but is restricted to fixed centerline connectivity and cannot flexibly accommodate topological changes such as varying branching patterns. In contrast, we propose a modular inference-time framework that controls unconditional diffusion models across diverse anatomical structures using geometric attributes such as size, shape, and position.
@@
Spatial Control of Generative Models. Spatial control of generative models is achieved through two principal approaches. The first approach conditions models on interpretable mid-level representations (e.g., bounding boxes, ellipsoid parameters, articulation angles) and has been successfully applied to images nie2024compositionalblobgen, video feng2025blobgen, 3D objects hertz2022spaghetti; koo2023salad; mu2021sdf, and proteins anonymous2024protcomposer. However, these methods cannot accommodate novel constraints without retraining. The second approach employs energy-based guidance during the reverse diffusion process bansal2023universal; du2023reduce, enabling flexible constraint composition at test time, but is typically limited to general localization rather than exact geometric control. Recent works such as self-guidance use attention-based loss functions to enable basic geometric attribute control (size, position) in text-to-image models epstein2023diffusionselfguidance, but is not designed for multi-label segmentations, nor does it control for orientation or aspect ratio. In our work, we extend energy-based guidance by introducing differentiable geometric losses for 3D multi-component anatomical voxel maps based on substructure-specific geometric properties such as the mass, centroid, and covariance, enabling the composition of multiple constraints across several independent substructures.
@@ 3 Anatomical Diffusion Models
Let $\mathbf{x}\in\mathbb{R}^{C\times H\times W\times D}$ be a 3D segmentation volume with $C$ tissue channels and $(H,W,D)$ spatial dimensions. We employ a variational autoencoder (VAE) with an encoder $\mathcal{E}$ that maps $\mathbf{x}$ to a lower-dimensional latent representation $\mathbf{z}=\mathcal{E}(\mathbf{x})$, and a decoder $\mathcal{D}$ that maps $\mathbf{z}$ back to voxel space reconstruction $\tilde{\mathbf{x}}=\mathcal{D}(\mathbf{z})$. The latent grid $\mathbf{z}\in\mathbb{R}^{c\times h\times w\times d}$ comprises $c$ channels and spatial dimensions (h,w,d)=(H/f,W/f,D/f)(h,w,d)=(H/f,\,W/f,\,D/f) for an integer downsampling factor $f$.
We use an unconditional latent diffusion model (LDM) as a prior over 3D anatomical segmentations, trained on the encoded latent representations $\mathbf{z}$. We specifically use the elucidated diffusion formulation of karras2022elucidating. In the forward process, data samples $\mathbf{z}\sim p_{\text{data}}(\mathbf{z})$ are progressively corrupted by adding Gaussian noise, yielding in perturbed data $\mathbf{z}_{\sigma}=\mathbf{z}+\mathbf{n}$ where $\mathbf{n}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$. The reverse process reconstructs the original data by approximating the score function $\nabla_{\mathbf{z}_{\sigma}}\log p(\mathbf{z}_{\sigma};\sigma)$ that controls the reverse diffusion process:
@@
$D_{\theta}(\mathbf{z}_{\sigma};\sigma)=c_{\text{skip}}(\sigma)\,\mathbf{z}_{\sigma}+c_{\text{out}}(\sigma)\,F_{\theta}(c_{\text{in}}(\sigma)\,\mathbf{z}_{\sigma};c_{\text{noise}}(\sigma))\,.$ (2)
@@ 4.2 Segmentation Denoising and Guidance
We formulate loss-based guidance in terms analogous to diffusion posterior sampling chung2022diffusionposteriorsampling, where the gradient derived from a differentiable geometric loss $\mathcal{L}_{\text{geom}}$ guides the sampling process. To guide anatomical generation, the intermediately noised latent $\mathbf{z}_{\sigma}$ is denoised by the diffusion model to produce a clean prediction $\mathbf{\hat{z}}_{0}=D_{\theta}(\mathbf{z}_{\sigma};\sigma)$ and subsequently decoded into a voxel-space segmentation $\mathbf{\hat{x}}_{0}=\mathcal{D}(\hat{\mathbf{z}}_{0})$. As the decoder outputs are continuous logits, we apply a label-wise softmax to ensure that the segmentation values are close to binary. The geometric loss $\mathcal{L}_{\text{geom}}$ is then computed in a differentiable manner to update the denoiser predictions through the gradient with respect to the noised latent $\mathbf{z}_{\sigma}$. The update step is parameterized with a guidance weight $w$ as follows:
@@ 4.3 Geometric Moment Loss
To isolate guidance to specific substructures representing individual tissues, we map the input segmentation $\hat{\mathbf{x}}_{0}\in\mathbb{R}^{C\times H\times W\times D}$ to a set of substructure voxel maps $\mathbf{\Omega}\in\mathbb{R}^{E\times H\times W\times D}$. Here, $E$ specifies the number of relevant substructures. Substructures are determined either through taking subsets of the tissue channels or taking the Boolean union of multiple tissue channels.
@@
where $\mathbf{1}^{T}$ is the all-ones vector, and $diag(\cdot)$ refers to diagonal matrix embedding. To enable independent control over size and shape characteristics, we compute a normalized representation of the covariance matrix. The scale-normalized covariance matrix is defined as $\mathcal{S}^{n}_{k}=\mathcal{S}_{k}/\operatorname{tr}(\Lambda)$ where $\Lambda$ is the eigenvalue matrix obtained from the eigendecomposition of $\mathcal{S}_{k}$. Intuitively, the normalized covariance matrix represents the aspect ratio and orientation of the substructure.
@@
Using prescribed weight factors $\lambda_{0},\lambda_{1},\lambda_{2}$, we compute the aggregate geometric loss as $\mathcal{L}_{\text{geom}}=\lambda_{0}\mathcal{L}_{\text{size}}+\lambda_{1}\mathcal{L}_{\text{pos}}+\lambda_{2}\mathcal{L}_{\text{shape}}\,$. The weighted sum of individual weights $\lambda_{i}$ allow us to control the contribution of individual loss to the guidance process, enabling easy disentangled control by zeroing out the associated weighting factor.
@@ 5.1 Unconditional Model Training
For diffusion training, we use the label maps provided in the TotalSegmentator dataset wasserthal2022totalsegmentator. We extract heart related labels, which include aorta (Ao), pulmonary artery (PA), pulmonary veins (PV), inferior vena cava (IVC), superior vena cava (SVC), left atrium (LA), right atrium (RA), left ventricle (LV), and right ventricle (RV), and left ventricular myocardium (Myo). We manually filter out low-quality label maps, resulting in 596 3D cardiac segmentations with 11 channels and an isotropic voxel edge length of $2\,\text{mm}$. See Section 8.2 for further details. We split the dataset into training and validation sets with an 80/20 split. All target moments and evaluation metrics are computed on the validation set.
We train an unconditional diffusion model on cardiac label maps similarly to kadry2024probing (further details in Section 8.3). To compute our geometric guidance loss, we use the weighted sum of the individual geometric moment losses, where the guidance weights $\lambda_{i}$ are tuned experimentally (see Section 8.4).
@@ 5.2 Baselines
We compare our approach (unconditional diffusion combined with geometric guidance) to conditional training approaches. Given the target geometric moments representing the size $\mathcal{M}$, centroid $\mathcal{C}$ and covariance $\mathcal{S}$ of each cardiac substructure, we condition the model in the following ways:
  • Explicit Concatenation: We directly encode geometric attributes as scalar values in the conditioning signal kadry2025diffusion. Here, we adapt this method to positional and shape-based features. We flatten and stack all geometric moments into a 13-dimensional vector for all $E$ substructures. We then expand this vector into a voxel grid $\mathcal{G}_{\text{exp}}\in\mathbb{R}^{13\times E\times h\times w\times d}$ which is concatenated to the latents along the channel dimension.
  • Implicit Concatenation: We indirectly encode geometric attributes in the conditioning signal through 3D heatmaps kadry2025diffusion. Here, we embed geometric moments as 3D Gaussians in voxel space. For each substructure, we create a voxel map $\mathcal{G}_{\text{imp}}\in\mathbb{R}^{E\times h\times w\times d}$ where the voxel values encode the Mahalanobis distance.
  • Cross-attention: We express the conditioning signal as a sequence of tokens where each token represents substructure geometry. The dimension of each token corresponds to the embedded geometric moments $\mathcal{G}_{\text{cross}}\in\mathbb{R}^{E\times 256}$. To enable sequence conditioning for the denoising U-Net, we convert the self-attention layers to cross-attention layers, similar to rombach2022ldm.
We implement guidance mechanisms such as adaptive null guidance (ANG) kadry2025diffusion for explicit concatenation, and classifier-free guidance (CFG) ho2022classifier for implicit concatenation and cross-attention. Further details can be found in Section 8.5.

### v3

@@ Abstract
Generative models of 3D cardiovascular anatomy can synthesize informative structures for clinical research and medical device evaluation, but face a trade-off between geometric controllability and realism. We propose CardioComposer: a programmable, inference-time framework for generating multi-class anatomical label maps from interpretable ellipsoidal primitives. These primitives represent geometric attributes such as the size, shape, and position of discrete substructures. We specifically develop differentiable measurement functions based on voxel-wise geometric moments, enabling loss-based gradient guidance during diffusion model sampling. We demonstrate that these losses can constrain individual geometric attributes in a disentangled manner and provide compositional control over multiple substructures. Finally, we show that our method is compatible with a broad range of anatomical systems containing non-convex substructures, spanning cardiac, vascular, and skeletal organs. We release our code at https://github.com/kkadry/CardioComposer.
@@ 1 Introduction
Three-dimensional segmentations of human anatomy power a variety of physics-based simulation platforms. For example, virtual cohorts of anatomy can be used for virtual clinical trials to evaluate and optimize novel medical devices and imaging systems (Sarrami-Foroushani et al., 2021; Viceconti et al., 2021; Abadi et al., 2020). Additionally, biophysical simulations can generate insights in the context of both computational physiology research (Niederer et al., 2020; Straughan et al., 2023; Roney et al., 2020) and surgical training (Yu et al., 2024). Anatomical segmentations can also be used to augment machine-learning workflows through the formation of synthetic images, either through imaging simulators (Gopalakrishnan et al., 2024; Gopalakrishnan and Golland, 2022), domain randomization (Dey et al., 2025; Billot et al., 2023), or generative models (Fernandez et al., 2022; 2024).
@@
Generative models of anatomy trained on patient-specific data offer advantages for simulation use-cases. For example, conditional generation can augment computational trial cohorts with anatomical variants that are both novel and rare (Kong et al., 2024). Moreover, generative editing methods, such as inpainting, can precisely modify existing patient geometries to create anatomically plausible variations (Kadry et al., 2024; 2025). These “digital siblings” can be used with biophysical simulators to model counterfactual scenarios that elucidate the relationship between anatomical form and function.
However, unlike generative modeling of 3D shapes for artistic purposes, generating anatomical models for biophysical simulations presents several unique challenges. The first concerns scale-critical features, in which minor geometric variations on the order of millimeters can cause major fluctuations in physiological behavior (Fabris et al., 2022; Sacco et al., 2018; Moore and Dasi, 2015). Second, anatomical structure exerts attribute-specific effects, in which geometric features such as size and position play different roles in determining biophysical outcomes (Kadry et al., 2021). Third, the geometry of multiple substructures interact in a compositional manner (Kadry et al., 2021; Bhalodia et al., 2018; Kong et al., 2024), where simulated outcomes are influenced by the collective arrangement of various substructures. Lastly, to interface with clinicians and device engineers, such generative models should be controllable via primitives that are interpretable and physiologically relevant.
@@
  • Differentiable Geometry for Anatomical Characterization: We introduce a set of differentiable geometric measurement functions that compute physiologically relevant anatomical features from a substructure label map. We specifically measure voxel-wise geometric moments, computing size via zeroth-order moments, position via first-order moments, and shape via scale-normalized second-order moments.
@@ 2 Background
Traditional Morphometric Modeling for Anatomy. Morphometry involves quantifying anatomical structure through geometric measurements. Traditional morphometric approaches measure intrinsic features such as length, area, volume, and shape, as well as extrinsic factors such as position and orientation. Geometric measurements enable applications such as cardiovascular risk stratification (Asheghan et al., 2023; Mahmod et al., 2024) and orthopedic diagnosis (Gatti et al., 2024). However, traditional morphometric approaches face two key challenges: they cannot represent complex relationships between features, and multiple distinct anatomies may map to the same high-level measurements. To address these limitations, we propose a framework in which an unconditional diffusion model is controlled by traditional morphometric features (size, shape, and position) to generate a variety of realistic anatomical structures, providing an approachable anatomical modeling interface for both clinical and engineering workflows involving numerical simulation.
@@
Generative Models for Numerical Simulation. Anatomical models in the form of 3D meshes or label maps serve as a crucial tool for studying form-function relationships through physical simulations, enabling both scientific discovery and medical device design. However, current approaches to create such models must trade off between fidelity and control. Simple geometric primitives, such as cylinders for coronary arteries (Dong et al., 2023) and truncated ellipsoids for cardiac chambers (Aróstica et al., 2025) offer parametric control but fail to capture anatomical realism. Data-driven approaches such as autoencoders (Dou et al., 2024; Qiao et al., 2025) represent anatomy in terms of global shape vectors, and can generate synthetic data for mechanistic studies of heart disease (Hermida et al., 2024; Williams et al., 2022). However, such approaches are limited in their ability to model interpretable geometric attributes. Deformation editing methods (Pham et al., 2023; 2024) allow for interpretable control of anatomical geometry, but are limited to modifying existing structures. Recently, diffusion-based approaches such as inpainting and partial diffusion have been used to edit patient-specific anatomy to create “digital siblings” (Kadry et al., 2024). However, such edits can induce undesirable morphological bias when applied to rare and pathological cases. To this end, recent studies have imposed anatomical features by explicitly providing scalar conditioning features during training. For example, de Wilde et al. (2025) trained a conditional model on thyroid segmentations, and  Kadry et al. (2025) introduced morpho-skeletal conditioning and guidance mechanisms for coronary arteries. However, both approaches rely on conditional training and are restricted to size-related variables such as volume or cross-sectional area. Similarly, Du et al. (2025) presented a hierarchical conditional diffusion model for generating aortic centerlines and radial profiles, but is restricted to fixed centerline connectivity and cannot flexibly accommodate topological changes such as varying branching patterns. In contrast, we propose a modular inference-time framework that controls unconditional diffusion models across diverse anatomical structures using geometric attributes such as size, shape, and position.
@@
Spatial Control of Generative Models. Spatial control of generative models is achieved through two principal approaches. The first approach conditions models on interpretable mid-level representations (e.g., bounding boxes, ellipsoid parameters, articulation angles) and has been successfully applied to images (Nie et al., 2024), videos (Feng et al., 2025), 3D objects (Hertz et al., 2022; Koo et al., 2023; Mu et al., 2021), and proteins (Stark et al., 2025). However, these methods cannot accommodate novel constraints without retraining. The second approach employs energy-based guidance during the reverse diffusion process (Bansal et al., 2023; Du et al., 2023), enabling flexible constraint composition at test time, but is typically limited to general localization rather than exact geometric control. Recent works such as self-guidance use attention-based loss functions to enable basic geometric attribute control (size, position) in text-to-image models (Epstein et al., 2023), but are not designed for multi-label segmentations, nor do they control for orientation or aspect ratio. In our work, we extend energy-based guidance by introducing differentiable geometric losses for 3D multi-component anatomical voxel maps based on substructure-specific geometric properties such as the mass, centroid, and covariance, enabling the composition of multiple constraints across several independent substructures.
@@ 3 Anatomical Diffusion Models
Let $\mathbf{x}\in\mathbb{R}^{C\times H\times W\times D}$ be a 3D segmentation volume with $C$ tissue channels and $(H,W,D)$ spatial dimensions. We employ a variational autoencoder (VAE) with an encoder $\mathcal{E}$ that maps $\mathbf{x}$ to a lower-dimensional latent representation $\mathbf{z}=\mathcal{E}(\mathbf{x})$, and a decoder $\mathcal{D}$ that maps $\mathbf{z}$ back to a voxel-space reconstruction $\tilde{\mathbf{x}}=\mathcal{D}(\mathbf{z})$. The latent grid $\mathbf{z}\in\mathbb{R}^{c\times h\times w\times d}$ comprises $c$ channels and spatial dimensions (h,w,d)=(H/f,W/f,D/f)(h,w,d)=(H/f,\,W/f,\,D/f) for an integer downsampling factor $f$.
We use an unconditional latent diffusion model (LDM) as a prior over 3D anatomical segmentations, trained on the encoded latent representations $\mathbf{z}$. We specifically use the elucidated diffusion formulation of Karras et al. (2022). In the forward process, data samples $\mathbf{z}\sim p_{\text{data}}(\mathbf{z})$ are progressively corrupted by adding Gaussian noise, resulting in perturbed data $\mathbf{z}_{\sigma}=\mathbf{z}+\mathbf{n}$ where $\mathbf{n}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$. The reverse process reconstructs the original data by approximating the score function $\nabla_{\mathbf{z}_{\sigma}}\log p(\mathbf{z}_{\sigma};\sigma)$ that controls the reverse diffusion process:
@@
$D_{\theta}(\mathbf{z}_{\sigma};\sigma)=c_{\text{skip}}(\sigma)\,\mathbf{z}_{\sigma}+c_{\text{out}}(\sigma)\,F_{\theta}(c_{\text{in}}(\sigma)\,\mathbf{z}_{\sigma};c_{\text{noise}}(\sigma))\,$, (2)
@@ 4.2 Segmentation Denoising and Guidance
We formulate loss-based guidance in terms analogous to diffusion posterior sampling (Chung et al., 2023), where the gradient derived from a differentiable geometric loss $\mathcal{L}_{\text{geom}}$ guides the sampling process. To guide anatomical generation, the intermediately noised latent $\mathbf{z}_{\sigma}$ is denoised by the diffusion model to produce a clean prediction $\mathbf{\hat{z}}_{0}=D_{\theta}(\mathbf{z}_{\sigma};\sigma)$ and subsequently decoded into a voxel-space segmentation $\mathbf{\hat{x}}_{0}=\mathcal{D}(\hat{\mathbf{z}}_{0})$. As the decoder outputs are continuous logits, we apply a label-wise softmax to ensure that the segmentation values are between 0 and 1. The geometric loss $\mathcal{L}_{\text{geom}}$ is then computed in a differentiable manner to update the denoiser predictions through the gradient with respect to the noised latent $\mathbf{z}_{\sigma}$. The update step is parameterized with a guidance weight $w$ as follows:
@@ 4.3 Geometric Moment Loss
To isolate guidance to specific substructures representing individual tissues, we map the input segmentation $\hat{\mathbf{x}}_{0}\in\mathbb{R}^{C\times H\times W\times D}$ to a set of substructure voxel maps $\mathbf{\Omega}\in\mathbb{R}^{E\times H\times W\times D}$. Here, $E$ specifies the number of relevant substructures. Substructures are determined either through taking subsets of the tissue channels or taking the union of multiple tissue channels.
@@
where $\mathbf{1}^{T}$ is the all-ones vector, and $diag(\cdot)$ refers to a diagonal matrix embedding. To enable independent control over size and shape characteristics, we compute a normalized representation of the covariance matrix. The scale-normalized covariance matrix is defined as $\mathcal{S}^{n}_{k}=\mathcal{S}_{k}/\operatorname{tr}(\Lambda)$ where $\Lambda$ is the eigenvalue matrix obtained from the eigendecomposition of $\mathcal{S}_{k}$. Intuitively, the normalized covariance matrix represents the aspect ratio and orientation of the substructure.
@@
Using prescribed weight factors $\lambda_{0},\lambda_{1},\lambda_{2}$, we compute the aggregate geometric loss as $\mathcal{L}_{\text{geom}}=\lambda_{0}\mathcal{L}_{\text{size}}+\lambda_{1}\mathcal{L}_{\text{pos}}+\lambda_{2}\mathcal{L}_{\text{shape}}\,$. The weighted sum of each weight $\lambda_{i}$ allows us to control the contribution of each individual loss to the guidance process, enabling easy disentangled control by zeroing out the associated weighting factor.
@@ 5.1 Unconditional Model Training
For diffusion training, we use the label maps provided in the TotalSegmentator dataset (Wasserthal et al., 2023). We extract heart-related labels, which include aorta (Ao), pulmonary artery (PA), pulmonary veins (PV), inferior vena cava (IVC), superior vena cava (SVC), left atrium (LA), right atrium (RA), left ventricle (LV), right ventricle (RV), and left ventricular myocardium (Myo). We manually filter out low-quality label maps, resulting in 596 3D cardiac segmentations with 11 channels and an isotropic voxel edge length of $2\,\text{mm}$ (Section 9.2). We split the dataset into training and validation sets with an 80/20 split. All target moments and evaluation metrics are computed on the validation set.
We train an unconditional diffusion model on cardiac label maps similarly to Kadry et al. (2024) (further details in Section 9.3). To compute our geometric guidance loss, we use the weighted sum of the individual geometric moment losses, where the guidance weights $\lambda_{i}$ are tuned experimentally (Section 9.4).
@@ 5.2 Baselines
We compare our approach (unconditional diffusion combined with geometric guidance) to conditional training approaches. Given the target geometric moments representing the size $\mathcal{M}$, centroid $\mathcal{C}$, and covariance $\mathcal{S}$ of each cardiac substructure, we condition the model in the following ways:
  • Explicit Concatenation: We directly encode geometric attributes as scalar values in the conditioning signal (Kadry et al., 2025). Here, we adapt this method to positional and shape-based features. We flatten and stack all geometric moments into a 13-dimensional vector for all $E$ substructures. We then expand this vector into a voxel grid $\mathcal{G}_{\text{exp}}\in\mathbb{R}^{(13\times E)\times h\times w\times d}$ which is concatenated to the latents along the channel dimension.
  • Implicit Concatenation: We indirectly encode geometric attributes in the conditioning signal through 3D heatmaps (Kadry et al., 2025). Here, we embed geometric moments as 3D Gaussians in voxel space. For each substructure, we create a voxel map $\mathcal{G}_{\text{imp}}\in\mathbb{R}^{E\times h\times w\times d}$ where the voxel values encode the Mahalanobis distance.
  • Cross-attention: We express the conditioning signal as a sequence of tokens where each token represents substructure geometry. The dimension of each token corresponds to the embedded geometric moments $\mathcal{G}_{\text{cross}}\in\mathbb{R}^{E\times 256}$. To enable sequence conditioning for the denoising U-Net, we convert the self-attention layers to cross-attention layers, similar to Rombach et al. (2022).
We implement guidance mechanisms such as adaptive null guidance (ANG) (Kadry et al., 2025) for explicit concatenation, and classifier-free guidance (CFG) (Ho and Salimans, 2022) for implicit concatenation and cross-attention. Further details can be found in Section 9.5.